# SpecCLIP: Aligning and Translating Spectroscopic Measurements for Stars

### v3

@@ Abstract
In recent years, large language models (LLMs) have transformed natural language understanding through vast datasets and large-scale parameterization. Inspired by this success, we present SpecCLIP, a foundation model framework that extends LLM-inspired methodologies to stellar spectral analysis. Stellar spectra, akin to structured language, encode rich physical and chemical information about stars. By training foundation models on large-scale spectral datasets, our goal is to learn robust and informative embeddings that support diverse downstream applications. As a proof of concept, SpecCLIP involves pre-training on two spectral types–LAMOST low-resolution and Gaia XP–followed by contrastive alignment using the CLIP (Contrastive Language-Image Pre-training) framework, adapted to associate spectra from different instruments. This alignment is complemented by auxiliary decoders that preserve spectrum-specific information and enable translation (prediction) between spectral types, with the former achieved by maximizing mutual information between embeddings and input spectra. The result is a cross-spectrum framework enabling intrinsic calibration and flexible applications across instruments. We demonstrate that fine-tuning these models on moderate-sized labeled datasets improves adaptability to tasks such as stellar-parameter estimation and chemical-abundance determination. SpecCLIP also enhances the accuracy and precision of parameter estimates benchmarked against external survey data. Additionally, its similarity search and cross-spectrum prediction capabilities offer potential for anomaly detection. Our results suggest that contrastively trained foundation models enriched with spectrum-aware decoders can advance precision stellar spectroscopy.
@@
1 Introduction
@@
Over the past decades, large-scale spectroscopic surveys have revolutionized our understanding of the formation and evolution of the Milky Way (Gilmore et al., 1989; Freeman & Bland-Hawthorn, 2002; Gray et al., 2002; Wyse, 2009; Helmi, 2020). These advances have been driven by three key forces. First, the continuous development of large-scale spectroscopic surveys – such as RAVE (Steinmetz et al., 2006), SEGUE (de Jong et al., 2010), APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015), LAMOST (Zhao et al., 2012), and DESI (DESI Collaboration et al., 2016) – has provided an unprecedented volume of stellar spectra across diverse Galactic populations. Second, the creation of powerful data infrastructures (Helou et al., 1991; Szalay & Gray, 2001; Gray et al., 2002; Fitzpatrick et al., 2014; Moitinho et al., 2017), exemplified by the SkyServer (Szalay et al., 2001) and CasJobs (OMullane et al., 2005) systems built for the Sloan Digital Sky Survey (SDSS, Margon, 1999; Abazajian et al., 2003, has democratized access to these datasets and enabled efficient large-scale analyses. Third, the refinement of algorithms for extracting physical parameters from these spectra has enabled increasingly precise stellar characterization.
@@
The latter effort encompasses traditional line-index methods, such as the SEGUE Stellar Parameter Pipeline (SSPP) (Lee et al., 2008); template-matching techniques, including UlySS (Koleva et al., 2009), the LAMOST stellar parameter pipeline (LASP; Wu et al., 2014) and the LAMOST stellar parameter pipeline at Peking University (LSP3; Xiang et al., 2015); as well as a range of machine learning approaches, among which the SSPP also incorporates a neural network module, alongside methods like the Cannon (Ness et al., 2015), the Payne (Ting et al., 2017, 2019), the DD-Payne (Xiang et al., 2019), and the TransformerPayne (Różański et al., 2025). All of these approaches, though diverse in methodology, rely heavily on supervision – either empirical or theoretical.
@@
Empirical approaches are limited by the coverage of the reference libraries they use, even when these libraries are derived from fundamental measurements. For instance, the official LAMOST stellar parameter pipeline (LASP; Wu et al., 2014), which is based on the UlySS algorithm, can only measure iron abundances down to [Fe/H] =2.5=-2.5, due to the limited parameter coverage of the ELODIE library (Moultaka et al., 2004). Theoretical approaches, while offering broader parameter coverage, are still subject to discrepancies between synthetic and observed spectra.
@@
Moreover, most existing pipelines are primarily designed to estimate atmospheric parameters, such as effective temperature ($T_{\rm eff}$), surface gravity (log $g$), and iron abundance ([Fe/H]), alongside a small set of elemental abundances. In contrast, determining other stellar properties, including reddening, stellar mass, and age, typically requires dedicated, task-specific pipelines. These efforts are further complicated by the diversity of spectral data in terms of wavelength coverage, resolution, and signal-to-noise ratio (SNR) – we refer to such quantities here as “modalities”. Achieving consistency and placing all inferred parameters onto a uniform scale remains a significant challenge in the prevailing framework, where each parameter is commonly derived using a distinct, often non-overlapping model.
@@
In parallel, the past five years have witnessed the remarkable success of large language models (LLMs) in natural language understanding, conversational AI, and text generation (Vaswani et al., 2017; Devlin et al., 2018; Radford et al., 2018, 2019; Brown et al., 2020). Breakthroughs in high-impact scientific domains – such as AlphaFold for protein structure prediction (Jumper et al., 2021) – have been enabled by the combination of massive datasets, large-scale models, and modern computational infrastructure (Kaplan et al., 2020).
@@
Stellar spectra can be analogized to a structured language: their rich absorption features and overall shapes encode key information about a star’s physical properties and evolutionary history. With the accumulation of millions of stellar spectra, it has become feasible to train foundation models (Leung & Bovy, 2024; Buck & Schwarz, 2024; Parker et al., 2024; Rizhko & Bloom, 2024; Smith et al., 2024; Zhong et al., 2024; Euclid Collaboration et al., 2025; Pattnaik et al., 2025) on these data using techniques inspired by LLMs. Here, “foundation models” refer to models pre-trained on large and diverse datasets. The broader the spectral distribution and parameter ranges used during pre-training, the better the model can learn the underlying structure of the “spectral language”. Once pre-trained, these models can be quickly fine-tuned with a small set of high-quality labels to perform a variety of downstream tasks, such as parameter estimation.
@@
A particularly promising framework for learning across modalities is the Contrastive Language–Image Pre-training (CLIP) algorithm (Radford et al., 2021), which aligns text and image representations via contrastive learning. CLIP jointly trains two encoders (one for each modality) by maximizing the similarity between representations of matched pairs and minimizing it for mismatched ones. The result is a shared embedding space that enables direct comparison between different modalities – allowing the model to retrieve or match one modality given the other, even for new inputs not seen during training.
@@
When applied to stellar spectra (Buck & Schwarz, 2024; Parker et al., 2024; Rizhko & Bloom, 2024), CLIP-style models can align spectra from different instruments or modalities with other astrophysical measurements, enabling more flexible downstream tasks such as parameter estimation and anomaly detection. However, a known limitation of CLIP is that it prioritizes the shared information between modalities, potentially discarding modality-specific features that are nonetheless important for some downstream tasks (Shwartz Ziv & LeCun, 2024).
@@
Partly motivated by this issue, we introduce the SpecCLIP project – a unified framework for cross-modal representation learning of stellar spectra. SpecCLIP begins with pre-training on two types of spectra: LAMOST (Cui et al., 2012) low-resolution spectra (LRS; Zhao et al., 2012) and Gaia XP spectra (De Angeli et al., 2023; Gaia Collaboration et al., 2023). These are aligned in a shared embedding space using a CLIP-like contrastive objective, enhanced by auxiliary decoders to preserve the mutual information between the learned embeddings and the input spectra.
@@
Mutual information (MI, Barber & Agakov, 2003; Poole et al., 2019; Devon Hjelm et al., 2018; Sui et al., 2023; Ting, 2025) [^1] Formally, the mutual information between two continuous random variables $X$ and $Y$ is defined as $I(X;Y)=\int\!\!\!\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}\,dx\,dy,$ where $p(x,y)$ is the joint density, and $p(x)$, $p(y)$ are the marginals. For discrete variables, the integral becomes a sum. is a key concept in information theory that quantifies how much one variable tells us about another. It serves as a natural objective for representation learning. While we do not explicitly compute MI between the spectra and their embeddings, we adopt a simple and effective strategy (Wang et al., 2022): augmenting the contrastive training with input reconstruction as a regularization mechanism – akin to using a decoder in an autoencoder to encourage informative representations.
@@
Another interesting feature enabled by our model is spectrum-to-spectrum translation. If the embeddings capture physically meaningful and shared information between spectra of different modalities, then it should be possible to predict one modality from the other, given suitable (spectra-to-spectra) supervision. In parallel, Buck & Schwarz (2024) demonstrated a similar idea using contrastive learning between Gaia XP and RVS spectra, employing CNN and multilayer perceptron (MLP) architectures. In our framework, we combine contrastive training and cross-modal prediction, augmented by spectral reconstruction within a unified architecture, enabling both spectrum-to-spectrum and spectrum-to-parameter applications.
@@
  • We propose extensions to the standard spectral CLIP model by incorporating auxiliary decoders alongside contrastive loss, encouraging the learning of both shared and non-shared spectral information.
@@
  • We demonstrate that our unified model enables both in-modal and cross-modal search, and cross-modal prediction.
@@
The remainder of this paper is organized as follows. In Section 2, we introduce the SpecCLIP model, including the separate pre-trained models and the CLIP-based alignment. Section 3 describes the downstream tasks, including modeling and sample selection for parameter estimation. Results, including parameter inference, spectral retrieval, and cross-modal prediction, are presented in Section 4. Section 5 offers further discussion, and Section 6 summarizes the work. Additional material is provided in the appendices: Appendix A (additional results); Appendix B (continuum fitting); Appendix C (normalizing flows for parameter inference); Appendix D (pre-training details); and Appendix E (projection models and decoders).
@@
Figure 1: Architecture of SpecCLIP. Two types of spectra (Shown here are examples of normalized LAMOST LRS and Gaia XP spectra varying with metalicities) are passed through two pre-trained spectral foundation models to obtain embeddings, where the pre-trained models can be either transformer-based networks or multilayer perceptron (MLP)-based autoencoders. These embeddings are then projected into a joint embedding space, which may optionally be split into a shared and a non-shared subspace. Based on the projected embeddings, we construct various loss functions to enable CLIP-like contrastive learning, cross-modal prediction, and spectral reconstruction. The combination of these loss functions results in five model variants: a baseline CLIP without decoders, CLIP-r with only reconstruction decoders, CLIP-p with only prediction decoders, CLIP-pr with full decoders, and CLIP-split with full decoders and an explicit separation of shared and non-shared embedding spaces (see Section 2.3)
2 SpecCLIP
@@
SpecCLIP, developed in this work, is a foundation model designed to align stellar spectra across different modalities, such as those with varying wavelength coverage, resolution, and signal-to-noise ratios, using a CLIP-inspired architecture. [^2] CLIP stands for Contrastive Language–Image Pre-training. In this work, we adopt the contrastive learning concept from the original CLIP framework, applying it to spectra–spectra pre-training. Although our use case differs from the original text–image alignment, we retain the term CLIP for consistency. Our approach builds upon transformer-based foundation models or MLP-based autoencoders tailored for specific spectral types. These are combined with contrastive learning to align different spectra using a standard contrastive loss, supplemented by additional modules to capture both shared and modality-specific information. Once trained, SpecCLIP enables various downstream tasks with either branch (not combination) of modalities, using a relatively small number of labeled examples (i.e., few-shot learning, as referred to in the literature).
@@
Figure 1 provides an overview of the SpecCLIP framework. Foundation models are first independently pre-trained on each spectral modality in an unsupervised fashion. These models are then aligned using contrastive learning, with auxiliary decoders added to enhance information retention. After this process, the final model is capable of handling a wide range of downstream tasks. Below, we describe the key experimental settings, including data-selection criteria, model architectures, loss functions for the foundation model training, and the connection of reconstruction loss to mutual information. Further implementation details are provided in Appendix D and Appendix E.
2.1 Pre-trained Foundation Model for LAMOST LRS
@@
The most recent LAMOST data release (DR11) [^3] https://www.lamost.org/dr11/ includes over 10 million low-resolution spectra (Luo et al., 2015), used for a variety of scientific tasks, including estimation of stellar parameters, chemical abundances, reddening, radial velocity, stellar mass, and age. Our aim in pre-training a foundation model for LAMOST LRS is to learn informative and transferable representations that support many of these tasks with a small to moderate amount of labeled data.
As a first step, we select a subset of 966,082 high-quality spectra for pre-training, using a 9:1 train–validation split (the same split ratio is used for the other modeling efforts described in the following two subsections). The selection criteria are: (1) signal-to-noise ratio (SNR) in the SDSS $g$-band greater than 50, and (2) apparent $g$-band magnitude less than 15.8. Additionally, we include all spectral types beyond the AFGK classes, while attempting to balance the four AFGK types themselves. We also aim to balance giant and dwarf stars; however, due to the intrinsic distribution of the dataset, the final ratio of dwarfs to giants is approximately 3.8:1. To focus on the most informative spectral features, we retain the 400–560 nm wavelength range, resulting in 1462 flux points per spectrum. The spectra are normalized before being input into the model; the normalization procedure is described in Appendix B.
@@
Each spectrum is tokenized into overlapping segments (tokens) of 20 flux points, with a span of 10 points, producing 146 tokens per spectrum. These are passed through 6 self-attention layers, each yielding a 768-dimensional token embedding. During training, 6 non-overlapping chunks (each 10444In previous versions, this number was misstated; the correct value used in the experiments is 10. tokens) are randomly masked to encourage robust representation learning. The resulting masked transformer (MT) model has a total of 42.7 million trainable parameters.
2.2 Pre-trained Foundation Model for Gaia XP Spectra

### v4

@@ Abstract
In recent years, large language models (LLMs) have transformed natural language understanding through vast datasets and large-scale parameterization. Inspired by this success, we present SpecCLIP, a foundation model framework that extends LLM-inspired methodologies to stellar spectral analysis. Stellar spectra, akin to structured language, encode rich physical and chemical information about stars. By training foundation models on large-scale spectral datasets, our goal is to learn robust and informative embeddings that support diverse downstream applications. As a proof of concept, SpecCLIP involves pre-training on two spectral types–LAMOST low-resolution and Gaia XP–followed by contrastive alignment using the CLIP (Contrastive Language-Image Pre-training) framework, adapted to associate spectra from different instruments. This alignment is complemented by auxiliary decoders that preserve spectrum-specific information and enable translation (prediction) between spectral types, the former being achieved by maximizing mutual information between embeddings and input spectra. The result is a cross-spectrum framework that enables intrinsic calibration and flexible applications across instruments. We demonstrate that fine-tuning these models on moderate-sized labeled datasets improves adaptability to tasks such as stellar-parameter estimation and chemical-abundance determination. SpecCLIP also enhances the accuracy and precision of parameter estimates bench-marked against external survey data. In addition, its similarity search and cross-spectrum prediction capabilities offer potential for anomaly detection. Our results suggest that contrastively trained foundation models enriched with spectrum-aware decoders can advance precision stellar spectroscopy. Our code SpecCLIP is publicly available on GitHub .
@@
software: SpecCLIP (zhao_2025_17824840), PyTorch (NEURIPS2019_9015), Astropy (2013A&A...558A..33A; 2018AJ....156..123A)
I Introduction
@@
Over the past decades, large-scale spectroscopic surveys have revolutionized our understanding of the formation and evolution of the Milky Way (1989ARA&A..27.. [^5] G; 2002ARA&A..40..487F; 10.1117/12.461524; 2009IAUS..258...11W; 2020ARA&A..58..205H). These advances have been driven by three key forces. First, the continuous development of large-scale spectroscopic surveys – such as RAVE (2006AJ....132.1645S), SEGUE (2010ApJ...714..663D), APOGEE (2017AJ....154...94M), GALAH (2015MNRAS.449.2604D), LAMOST (2012RAA....12..723Z), and DESI (2016arXiv161100036D) – has provided an unprecedented volume of stellar spectra across diverse Galactic populations. Second, the creation of powerful data infrastructures (1991ASSL..171...89H; doi:10.1126/science.293.5537.2037; 10.1117/12.461524; 10.1117/12.2057445; 2017A&A...605A..52M), exemplified by the SkyServer (2001cs.......11015S) and CasJobs (2005cs........2072O) systems built for the Sloan Digital Sky Survey (SDSS, 1999RSPTA.357...93M; 2003AJ....126.2081A, has democratized access to these datasets and enabled efficient large-scale analyses. Third, the refinement of algorithms for extracting physical parameters from these spectra has enabled increasingly precise stellar characterization.
@@
The latter effort encompasses traditional line-index methods, such as the SEGUE Stellar Parameter Pipeline (SSPP) (2008AJ....136.2022L); template-matching techniques, including UlySS (2009A&A...501.1269K), the LAMOST stellar parameter pipeline (LASP; 2014IAUS..306..340W) and the LAMOST stellar parameter pipeline at Peking University (LSP3; 2015MNRAS.448..822X); as well as a range of machine learning approaches, among which the SSPP also incorporates a neural network module, alongside methods such as the Cannon (2015ApJ...808...16N), the Payne (2017ApJ...849L...9T; 2019ApJ...879...69T),the DD-Payne (2019ApJS..245...34X), and the TransformerPayne (2025ApJ...980...66R). All of these approaches, though diverse in methodology, rely heavily on supervision – either empirical or theoretical.
@@
Empirical approaches are limited by the coverage of the reference libraries they use, even when these libraries are derived from fundamental measurements. For instance, the official LAMOST stellar parameter pipeline (LASP; 2014IAUS..306..340W), which is based on the UlySS algorithm, can only measure iron abundances down to [Fe/H] =2.5=-2.5, due to the limited parameter coverage of the ELODIE library (2004PASP..116..693M). Theoretical approaches, while offering broader parameter coverage, are still subject to discrepancies between synthetic and observed spectra.
@@
In addition, most existing pipelines are designed primarily to estimate atmospheric parameters, such as effective temperature ($T_{\rm eff}$), surface gravity (log $g$) and iron abundance ([Fe/H]), in conjunction with a small set of elemental abundances. In contrast, determining other stellar properties, including reddening, stellar mass, and age, typically requires dedicated, task-specific pipelines. These efforts are further complicated by the diversity of spectral data in terms of wavelength coverage, resolution, and signal-to-noise ratio (SNR) – we refer to such quantities here as “modalities”. Achieving consistency and placing all inferred parameters onto a uniform scale remains a significant challenge in the prevailing framework, where each parameter is commonly derived using a distinct, often non-overlapping model.
@@
In parallel, the past five years have witnessed the remarkable success of large language models (LLMs) in natural language understanding, conversational AI, and text generation (NIPS2017_3f5ee243; 2018arXiv181004805D; radford2018improving; radford_language_2019; NEURIPS2020_1457c0d6). Breakthroughs in high-impact scientific domains – such as AlphaFold for protein structure prediction (jumper2021highly) – have been enabled by the combination of massive datasets, large-scale models, and modern computational infrastructure (journals/corr/abs-2001-08361).
@@
Stellar spectra can be analogized to a structured language: their rich absorption features and overall shapes encode key information about a star’s physical properties and evolutionary history. With the accumulation of millions of stellar spectra, it has become feasible to train foundation models (2024MNRAS.527.1494L; 2024arXiv241016081B; 2024MNRAS.531.4990P; 2024arXiv241108842R; 2024arXiv240514930S; 2024arXiv241221130Z; 2025arXiv250315312E; 2025arXiv250101070P) on these data using techniques inspired by LLMs. Here, “foundation models” refer to models pre-trained on large and diverse datasets. The broader the spectral distribution and parameter ranges used during pre-training, the better the model can learn the underlying structure of the “spectral language”. Once pre-trained, these models can be quickly fine-tuned with a small set of high-quality labels to perform a variety of downstream tasks, such as parameter estimation.
@@
A particularly promising framework for learning across modalities is the Contrastive Language–Image Pre-training (CLIP) algorithm (2021arXiv210300020R), which aligns text and image representations via contrastive learning. CLIP jointly trains two encoders (one for each modality) by maximizing the similarity between representations of matched pairs and minimizing it for mismatched ones. The result is a shared embedding space that enables direct comparison between different modalities, allowing the model to retrieve or match one modality given the other, even for new inputs not seen during training.
@@
When applied to stellar spectra (2024arXiv241016081B; 2024MNRAS.531.4990P; 2024arXiv241108842R), CLIP-style models can align spectra from different instruments or modalities with other astrophysical measurements, allowing for more flexible downstream tasks such as parameter estimation and anomaly detection. However, a known limitation of CLIP is that it prioritizes the shared information between modalities, potentially removing modality-specific features that are still important for some downstream tasks (2024Entrp..26..252S).
@@
Partly motivated by this issue, we introduce the SpecCLIP project – a unified framework for cross-modal representation learning of stellar spectra. SpecCLIP begins with pre-training on two types of spectra: LAMOST (2012RAA....12.1197C) low-resolution spectra (LRS; 2012RAA....12..723Z) and Gaia XP spectra (2023A&A...674A...2D; 2023A&A...674A...1G). These are aligned in a shared embedding space using a CLIP-like contrastive objective, enhanced by auxiliary decoders to preserve the mutual information between the learned embeddings and the input spectra.
@@
Mutual information (MI, 10.5555/2981345.2981371; 2019arXiv190506922P; 2018arXiv180806670D; 2023mla..confE..30S; 2025arXiv250612230T) [^1] Formally, the mutual information between two continuous random variables $X$ and $Y$ is defined as $I(X;Y)=\int\!\!\!\int p(x,y)\log\frac{p(x,y)}{p(x)p(y)}\,dx\,dy,$ where $p(x,y)$ is the joint density, and $p(x)$, $p(y)$ are the marginals. For discrete variables, the integral becomes a sum. is a key concept in information theory that quantifies how much one variable tells us about another. It serves as a natural objective for representation learning. Although we do not explicitly compute MI between the spectra and their embeddings, we adopt a simple and effective strategy (2022arXiv220307004W): increasing contrastive training with input reconstruction as a regularization mechanism, similar to using a decoder in an autoencoder to encourage informative representations.
@@
Another interesting feature enabled by our model is spectrum-to-spectrum translation. If the embeddings capture physically meaningful and shared information between spectra of different modalities, then it should be possible to predict one modality from the other, given suitable (spectra-to-spectra) supervision. In parallel, 2024arXiv241016081B demonstrated a similar idea using contrastive learning between Gaia XP and RVS spectra, employing CNN and multilayer perceptron (MLP) architectures. In our framework, we combine contrastive training and cross-modal prediction, augmented by spectral reconstruction within a unified architecture, enabling both spectrum-to-spectrum and spectrum-to-parameter applications.
@@
  • We demonstrate that our unified model enables both in-modal and cross-modal search, as well as cross-modal prediction.
@@
The remainder of this paper is organized as follows. In Section II, we introduce the SpecCLIP model, including the separate pre-trained models and the CLIP-based alignment. Section III describes the downstream tasks, including modeling and sample selection for parameter estimation. The results, including parameter inference, spectral retrieval, and cross-modal prediction, are presented in Section V. Section VI offers further discussion and Section VII summarizes the work. Additional material is provided in the appendices: Appendix A (additional results); Appendix B (continuum fitting); Appendix C (normalizing flows for parameter inference); Appendix D (pre-training details); Appendix E (projection models and decoders); and Appendix G (summary of key hyper-parameters and configurations)
@@
Figure 1: Architecture of SpecCLIP. Two types of spectra (Shown here are examples of normalized LAMOST LRS and Gaia XP spectra varying with metallicities) are passed through two pre-trained spectral foundation models to obtain embeddings, where the pre-trained models can be either transformer-based networks or multilayer perceptron (MLP)-based autoencoders. These embeddings are then projected into a joint embedding space, which may optionally be split into a shared and a non-shared subspace. Based on the projected embeddings, we construct various loss functions to enable CLIP-like contrastive learning, cross-modal prediction, and spectral reconstruction. The combination of these loss functions results in five model variants: a baseline CLIP without decoders, CLIP-r with only reconstruction decoders, CLIP-p with only prediction decoders, CLIP-pr with full decoders, and CLIP-split with full decoders and an explicit separation of shared and non-shared embedding spaces (see Section II.3)
II SpecCLIP
@@
SpecCLIP, developed in this work, is a foundation model designed to align stellar spectra across different modalities, such as those with varying wavelength coverage, resolution, and SNRs, using a CLIP-inspired architecture. [^2] CLIP stands for Contrastive Language–Image Pre-training. In this work, we adopt the contrastive learning concept from the original CLIP framework, applying it to spectra–spectra pre-training. Although our use case differs from the original text–image alignment, we retain the term CLIP for consistency. Our approach builds upon transformer-based foundation models or MLP-based autoencoders tailored for specific spectral types. These are combined with contrastive learning to align different spectra using a standard contrastive loss, supplemented by additional modules to capture both shared and modality-specific information. Once trained, SpecCLIP enables various downstream tasks with either branch (not combination) of modalities, using a relatively small number of labeled examples (i.e., few-shot learning, as referred to in the literature).
@@
Figure 1 provides an overview of the SpecCLIP framework. Foundation models are first independently pre-trained on each spectral modality in an unsupervised fashion. These models are then aligned using contrastive learning, with additional decoders added to enhance information retention. After this process, the final model is capable of handling a wide range of downstream tasks. Below, we describe the key experimental settings, including data-selection criteria, model architectures, loss functions for the foundation model training, and the connection of reconstruction loss to mutual information. Further details of the implementation are provided in Appendix D and Appendix E.
II.1 Pre-trained Foundation Model for LAMOST LRS
@@
The most recent LAMOST data release (DR11) [^3] https://www.lamost.org/dr11/ includes more than 10 million low-resolution spectra (2015RAA....15.1095L), used for a variety of scientific tasks, including estimation of stellar parameters, chemical abundances, reddening, radial velocity, stellar mass, and age. Our aim in pre-training a foundation model for LAMOST LRS is to learn informative and transferable representations that support many of these tasks with a small to moderate amount of labeled data.
@@
In our pre-trained LRS model, each spectrum is tokenized into overlapping segments (tokens) of 20 flux points 20 flux points ($\approx 22$ Å at a wavelength of 50005000Å, corresponding to $\sim$10 resolution elements and capturing typical local line structures), with the overlap of 10 points to preserve continuity, producing 146 tokens per spectrum. These tokens, together with one special token–the logarithmic standard deviation of the spectrum–are passed through 6 self-attention layers, each yielding a 768-dimensional token embedding. During training, 6 non-overlapping chunks (each 10 tokens, corresponding to an overall effective masking rate of $\approx$ 45%) are randomly masked to encourage robust representation learning. The resulting transformer-like framework with mask modeling (denoted as the masked transformer, or MT) has a total of 42.7 million trainable parameters.
II.2 Pre-trained Foundation Model for Gaia XP Spectra